OpenStudy (anonymous):

the half life of carbon 14 is 5730 years. How many grams of parent material will remain from a 100 gram sample of carbon 14 after 2 half lives

OpenStudy (anonymous):

\[N=N_0 e^{-kt}\] where n0 = starting mass n = mass after t years k = rate t = years is this formula right?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

if it is, then plug the values for half life to find the rate \[\huge 1 = 2e^{-k5730}\] find k then substitute it in the formula with t = half life time x 2 (5730 x 2) and n0 = 100

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

kinda?

OpenStudy (anonymous):

k = 0.000120968 \[\huge N = 100e^{0.000120968 \times 11460}\]~

OpenStudy (lgbasallote):

are you sure about this @nphuongsun93 ?

OpenStudy (anonymous):

Hope it's OK to butt in here. Since the problem gives an integral number of half-life(s), the problem is a lot easier. The amount remaining after 2 of them is just 1/2 times 1/2 = 1/4.

OpenStudy (anonymous):

yea... actually not too sure now D: after two half life would just equal 1/4 of the original mass.

OpenStudy (anonymous):

oh nvm my way got the same answer anyway l0l

OpenStudy (anonymous):

Also, given the half-life, the rate k times the half life T = ln 2

OpenStudy (anonymous):

OK. Good.